Minimize Q=1159+2L^0.5+5K^0.5 using Lagrangian method and constraint 1200=20L+30K
1
Expert's answer
2020-09-15T10:10:27-0400
"Solution"

maximizing

"Q=f(L,K)=1159+2L^{0.5}+5K^{0.5}"

Subject to the constraint

"20L+30K=1200"

Step 1: Create a new Lagrangian equation from the original information;

The Lagrangian is



"\theta=f(L,K)=1159+2L^{0.5}+5K^{0.5}+\lambda (1200-20L-30K)"

Step 2: Then follow the same steps as used in regular minimization problem.


The first order conditions are;

"\frac{\delta \theta}{\delta L}=\theta_L=l^{-0.5}-20\lambda=0\\
\frac{\delta \theta}{\delta K}=\theta_K=2.5K^{-0.5}-30\lambda=0\\
\frac{\delta \theta}{\delta \lambda}=1200-20L-30K=0\\"

From the first two equations we get


"3L^{-0.5}=5K^{-0.5}\\
5L^{0.5}=3K^{0.5}\\\\
L=\frac{\sqrt{15}}{5}K"

Substitute this result into the third equation


"1200-20(\frac{\sqrt{15}}{5})-30K=0\\
K=3.484"

Therefore


"L=2.699,\ \lambda=0.03"

Therefore, the combination "2.699\ units" of labor and "3.383\ units" of capital minimize the total cost of producing "1,200\ units" . In addition, "\u03bb" equals "0.03" .If the firm wants to produce one more unit of the good, the total cost increases by $0.03.